# Discrete Duality Symmetry in a $3D$ Field-Theoretic Model

### v2

@@ 2 Preliminaries: Constraint Analysis, (Dual-)Gauge and Discrete Duality Symmetry Transformations
The above equation demonstrates that if we choose the infinitesimal dual-gauge transformation parameters such that the restrictions: $\Box\Sigma_{\mu}=0$ and $\Box\Sigma=0$ are satisfied together, we have a perfect dual-gauge symmetry in the theory because the gauge-fixed Lagrangian density transforms to a total spacetime derivative (under the above conditions). In addition to the infinitesimal (i) gauge symmetry transformations (2), and (ii) dual-gauge symmetry transformations (14) (with similar kinds of restrictions on the transformation parameters: $\Box\Lambda=0,\;\Box\Lambda_{\mu}=0,\;\Box\Sigma=0,\;\Box\Sigma_{\mu}=0$), we have the following perfect discrete duality symmetry transformations§§§Mathematical origin for the existence of the perfect duality symmetry transformations (16) for the $3D$ combined system of the free Abelian 1-form and 2-form gauge theories, seems to be hidden in the Hodge duality * operation on the Abelian 2-form and 1-form, respectively. For instance, we note that: $*\,B^{(2)}=\big{[}\frac{1}{2}\,\varepsilon_{\mu\nu\sigma}\,B^{\nu\sigma}\big{]}\,dx^{\mu}\sim A_{\mu}\,dx^{\mu}$ which demonstrate that, for our $3D$ theory, the Hodge dual of 2-form is the 1-form. In a similar manner, we observe that the Hodge dual of the 1-form: $*\,A^{(1)}=\frac{1}{2}\,\big{[}\varepsilon_{\mu\nu\sigma}\,A^{\sigma}\big{]}\,(dx^{\mu}\wedge dx^{\nu})\sim\frac{1}{2}\,B_{\mu\nu}\,(dx^{\mu}\wedge dx^{\nu})$ is the 2-form. These observations are the reasons behind the numeral factors that appear in the discrete duality transformations (16) modulo the overall factors of ±$\pm\,i$.:
@@
where (i) the gauge-fixing term of the Abelian 1-form field interchanges with the kinetic term of the Abelian 2-form field, and (ii) the gauge-fixing term of the Abelian 2-form field interchanges with the kinetic term of the Abelian 1-form field. This observation is very crucial as it will play a key role at the quantum level when we shall discuss various kinds of continuous and discrete symmetries of our $3D$ system (as the BRST-quantized theory).
@@ 3 BRST and co-BRST Invariant Lagrangian Density
The gauge-fixed Lagrangian density (12) for our $3D$ combined system of the free Abelian 1-form and 2-form gauge theories can be linearized, in its most general form in two different ways, as follows
(,${\cal L}^{(1)}_{({\cal B},B)}$ $=$ $\dfrac{1}{2}\,{\cal B}^{\mu}{\cal B}_{\mu}-{\cal B}^{\mu}\Big{(}\varepsilon_{\mu\nu\sigma}\partial^{\nu}A^{\sigma}-\dfrac{1}{2}\,\partial_{\mu}\widetilde{\phi}\Big{)}-B\,(\partial\cdot A)+\dfrac{B^{2}}{2}$ (18)
@@
where the subscripts on the above Lagrangian densities denote the short forms of the Nakanishi-Lautrup auxiliary fields that have been invoked to linearize the kinetic and gauge-fixing terms of the combined system of the free Abelian 2-form and 1-form gauge theories. To be more specific, we note that, in the Lagrangian density (18), a pair of auxiliary fields (${\cal B}_{\mu},\,B$) have been invoked to linearize the kinetic and gauge-fixing terms of the $3D$ Abelian 1-form gauge theory, respectively. On the other hand, another pair of auxiliary fields (${\cal B},\,B_{\mu}$) have been incorporated into the theory to linearize the kinetic and gauge-fixing terms of the 3D Abelian 2-form theory, respectively. In exactly similar fashion, we can specify the the concise form of the subscript on the Lagrangian density (19) where the replacements for the pairs of auxiliary fields are: $({\cal B}_{\mu},\,B)\Rightarrow(\bar{\cal B}_{\mu},\,B)$ and $({\cal B},\,B_{\mu})\Rightarrow({\cal B},\,\bar{B}_{\mu})$, respectively. In the above Lagrangian densities (18) and (19), an additional pair of dynamical fields ($\widetilde{\phi},\;\phi$) have been added at appropriate places with proper mass dimensions in natural units (where $\hbar=c=1$). These fields are nothing but the pseudo-scalar and scalar fields, respectively. The above linearizations are general in the sense that we have taken into account both the signs that can be associated with the fields ($\widetilde{\phi},\;\phi$). In other words, a close look at (18) and (19) demonstrates that both the Lagrangian densities can be exchanged with each-other by taking into account the replacements: $(\widetilde{\phi},\;\phi)\leftrightarrow(-\,\widetilde{\phi},\;-\,\phi),\;({\cal B}_{\mu},\,B)\leftrightarrow(\bar{\cal B}_{\mu},\,B),\;({\cal B},\,B_{\mu})\leftrightarrow({\cal B},\,\bar{B}_{\mu})$. At this juncture, it is interesting to point out that the discrete duality symmetry transformations (16), existing for the gauge-fixed Lagrangian density (12), can be generalized into the following transformations:
@@
For the sake of brevity, in our present endeavor, we consider only (i) the off-shell nilpotent (co-)BRST symmetry transformations which are the generalizations of the classical (dual-)gauge symmetry transformations (14) and (2) to the quantum level, and (ii) the (co-)BRST invariant Lagrangian density ${\cal L}_{B}$ which is the generalization of the classical linearized version of the Lagrangian density (,${\cal L}^{(1)}_{({\cal B},B)}$ [cf. Eq. (18)] to its quantum level. In other words, we promote the Lagrangian density (18) to incorporate into itself the Faddeev-Popov (FP) ghost terms corresponding to the $3D$ field-theoretic model of the combined system of the Abelian 1-form and 2-form gauge theories as follows
@@
where the FP-terms, corresponding to the free massless Abelian 2-form theory, have been systematically derived in our earlier work [5]. In this work, we have considered the BRST-quantized version of the $4D$ Stückelberg-modified massive Abelian 2-form theory and the FP-ghost term (i.e. $-\,\partial^{\mu}\bar{C}\,\partial_{\mu}C$) for the free Abelian 1-form theory is the standard book material. In the above Lagrangian density (21), the Lorentz vector fermionic (i.e. $C_{\mu}^{2}=\bar{C}_{\mu}^{2}=0,\,C_{\mu}\,\bar{C}_{\nu}+\bar{C}_{\nu}\,C_{\mu}=0,\,C_{\mu}\,C_{\nu}+C_{\nu}\,C_{\mu}=0,$ etc.) (anti-)ghost fields $(\bar{C}_{\mu})C_{\mu}$ are the generalizations of the gauge transformations parameter $\Lambda_{\mu}$ [cf. Eq. (2)] and they carry the ghost numbers $(-1)+1$, respectively, and the bosonic (anti-)ghost fields $(\bar{\beta})\beta$ are endowed with the ghost numbers $(-2)+2$. Another pair of the fermionic (i.e. $C^{2}=\bar{C}^{2}=0,\,C\,\bar{C}+\bar{C}\,C=0$) (anti-)ghost fields $(\bar{C})C$ carry the ghost numbers $(-1)+1$ and they are the generalizations of the Lorentz scalar gauge transformation parameter $\Lambda$ [cf. Eq. (2)] and the auxiliary (anti-)ghost fields $(\rho)\lambda$ are endowed with ghost numbers $(-1)+1$, respectively. These (anti-)ghost fields are required to maintain the sacrosanct property of unitarity at any arbitrary order of perturbative computations. The other symbols in (21) carry the same meaning as explained in the context of the Lagrangian density in (18).
@@
$C_{\mu}\,\longrightarrow\,\pm\,i\,\bar{C}_{\mu},\qquad\quad\bar{C}_{\mu}\,\longrightarrow\,\pm\,i\,C_{\mu},\qquad\quad C\,\longrightarrow\,\mp\,i\,\bar{C},$
$\bar{C}\,\longrightarrow\,\mp\,i\,C,\qquad\quad\quad\beta\,\longrightarrow\,\pm\,i\,\bar{\beta},\qquad\quad\quad\bar{\beta}\,\longrightarrow\,\mp\,i\,\beta,$
$\rho\,\longrightarrow\,\mp\,i\,\lambda,~{}\qquad\quad\quad\lambda\,\longrightarrow\,\mp\,i\,\rho.$ (25)
Thus, it is clear that the total (co-)BRST invariant Lagrangian density (21) respects the discrete duality symmetry transformations in the non-ghost sector [cf. Eq. (3)] as well as in the ghost-sector [cf. Eq. (3)]. To sum-up, the Lagrangian density (21) respects the infinitesimal, continuous and nilpotent symmetries [cf. Eqs. (3),(3)] and a set of discrete duality symmetry transformations. The latter is the sum of the discrete symmetry transformations that have been listed in (3) and (3)). We call these discrete symmetry transformations as the duality symmetry transformations because the basic 1-form and 2-form fields are connected with each-other [cf. Eq. (3)] and they are dual to each-other on a flat $3D$ Minkowskian spacetime manifold [cf. footnote below Eq. (16)].
@@ 4 Bosonic Symmetry and Algebraic Structures
It is straightforward to note, from the equations (3) and (3), that the anticommutator (i.e. $\{s_{b},\;s_{d}\}$) between them is not equal to zero. In fact, this anticommutator defines a bosonic symmetry (i.e. $s_{\omega}=\{s_{b},\;s_{d}\}$), under which, the fields of (21) transform as:
$s_{\omega}B_{\mu\nu}=\varepsilon_{\mu\nu\sigma}\,\partial^{\sigma}B+\big{(}\partial_{\mu}{\cal B}_{\nu}-\partial_{\nu}{\cal B}_{\mu}\big{)},\qquad s_{\omega}C_{\mu}=\partial_{\mu}\lambda,$
$s_{\omega}A_{\mu}=-\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}B^{\sigma}+\partial_{\mu}{\cal B},\qquad s_{\omega}\bar{C}_{\mu}=\partial_{\mu}\rho,$
$s_{\omega}\big{[}B_{\mu},\,{\cal B}_{\mu},\,B,\,{\cal B},\,\phi,\,\widetilde{\phi},\,\beta,\,\bar{\beta},\rho,\,\lambda,\,C,\,\bar{C}\big{]}=0.$ (26)
At this stage, is worthwhile to mention that under the above bosonic symmetry transformations, the (anti-)ghost fields either do not transform at all or they transform up to the $U(1)$ gauge symmetry-type transformations. Furthermore, it is very interesting to point out that, in their operator forms, the nilpotent (co-)BRST symmetry transformations and the bosonic symmetry transformation obey the following algebra, namely;
@@
which establishes the fact that the bosonic symmetry transformation, in its operator form, commutes with the off-shell nilpotent (co-)BRST symmetry transformation operators. The algebra (4) is reminiscent of the following algebra obeyed by a set of three de Rham cohomological operators of differential geometry [14-18]
@@
where $d$ (with $d^{2}=0$) is the exterior derivative, $\delta=\pm\,*\,d\,*$ (with $\delta^{2}=0$) is the co-exterior (or dual-exterior) derivative and $\Delta=(d+\delta)^{2}$ is the Laplacian operator (which is always positive definite). Here the symbol * stands for the Hodge duality operator on the compact spacetime manifold on which the cohomological operators are defined.
We end this short section with a couple of useful remarks. First of all, at the algebraic level, it is very tempting to appreciate the similarity between the above algebraic relationships in (4) and (4) which allow us to state that there is a one-to-one mappingWe have obtained the one-to-one mapping because we have considered only the Lagrangian density (21) which is the generalization of the linearized classical version of the gauge-fixed Lagrangian density (18) to its quantum level. If we had considered the quantum version of the linearized classical version of the gauge-fixed Lagrangian density (19) together with the above generalization, we would have ended up with the two-to-one mapping between the symmetry transformation operators and the de Rham cohomological operators of differential geometry as we have obtained in our earlier works [6-9]. between the symmetry transformation operators ($s_{b},\;s_{d},\;s_{\omega}$) and the de Rham cohomological operators ($d,\;\delta,\;\Delta$). To corroborate this claim, we have to recall our key observations in Secs. 2 and 3 where we have established that, under the (dual-)gauge and (co-)BRST symmetry transformations, the (i) gauge-fixing terms (owing their origin to the co-exterior derivative), and (ii) kinetic terms (owing their existence to the exterior derivative) for the gauge fields remain invariant, respectively. Hence, the above identification: $(s_{b},\;s_{d},\;s_{\omega})\Leftrightarrow(d,\;\delta,\;\Delta)$ is realistic at the algebraic level. Second, the (co-)exterior derivatives $(\delta)d$, as pointed out earlier, are connected by the relationship: $\delta=\pm\,*\,d\,*$. Within the framework of BRST formalism, this relationship is realized in terms of the nilpotent (co-)BRST transformation operators and the discrete transformation operators [cf. Eqs. (3),(3)] as
@@
where the symbol * stands for the discrete duality symmetry transformations. Thus, we observe that the interplay between the continuous and discrete symmetry operators of our theory provide the physical realization of the relationship: $\delta=\pm\,*\,d\,*$. In the above equation (29), as is obvious, the generic field of the Lagrangian density (21) has been denoted by the field $\Phi$. The ($\pm$) sign, on the r.h.s. of the above equation (29), is determined and dictated by a couple of successive operations of the discrete duality symmetry transformation operator on the generic field $\Phi$ of the Lagrangian density (21) as follows [31]:
@@
It is an elementary exercise to note that, for all the bosonic fields of our (co-)BRST invariant Lagrangian density (21), we have the plus sign on the r.h.s. of (30) and, hence, in the equation (29), too. For instance, it is straightforward to check that: $*\,(*\,B_{\mu\nu})=+\,B_{\mu\nu}$. As a consequence, we observe that: $s_{d}B_{\mu\nu}=+\,*\,s_{b}\,*\,B_{\mu\nu}$ is true. On the other hand, the signs on the r.h.s. of (30) and (29) are negative when take into account the fermionic fields of the (co-)BRST invariant Lagrangian density (21). For instance, we can readily check that: $*\,(*\,C_{\mu})=-\,C_{\mu}$. As a consequence, we have the algebraic relationship between the off-shell nilpotent (co-)BRST symmetry transformations [cf. Eqs. (3),(3)] as: $s_{d}C_{\mu}=+\,*\,s_{b}\,*\,C_{\mu}$ where the symbol * stands for the set of discrete duality symmetry transformation operators that have been listed in (3) and (3).

### v4

@@ 2 Preliminaries: Constraint Analysis, (Dual-)Gauge and Discrete Duality Symmetry Transformations
The above equation demonstrates that if we choose the infinitesimal dual-gauge transformation parameters such that the restrictions: $\Box\Sigma_{\mu}=0$ and $\Box\Sigma=0$ are satisfied together, we have a perfect dual-gauge symmetry invariance in the theory because the gauge-fixed Lagrangian density (12) transforms to a total spacetime derivative (under the above conditions). To be precise, we note that, under the above restrictions on the transformation parameters, the action integral (corresponding to the gauge-fixed Lagrangian density (12)) remains invariant because of the Gauss divergence theorem due to which all the physical fields as well as the transformation parameters of our theory vanish off as $x\to\pm\,\infty$. In addition to the infinitesimal (i) gauge symmetry transformations (2), and (ii) dual-gauge symmetry transformations (14) (with similar kinds of restrictions on the transformation parameters: $\Box\Lambda=0,\;\Box\Lambda_{\mu}=0,\;\Box\Sigma=0,\;\Box\Sigma_{\mu}=0$), we have the following perfect discrete duality symmetry transformationsMathematical origin for the existence of the perfect duality symmetry transformations (16) for the $3D$ combined system of the free Abelian 1-form and 2-form gauge theories, seems to be hidden in the Hodge duality * operation on the Abelian 2-form and 1-form, respectively. For instance, we note that: $B^{(2)}=\big{[}\frac{1}{2!}\,\varepsilon_{\mu\nu\sigma}\,B^{\nu\sigma}\big{]}\,dx^{\mu}\sim A_{\mu}\,dx^{\mu}$ which demonstrate that, for our $3D$ theory, the Hodge dual of 2-form is the 1-form. In a similar manner, we observe that the Hodge dual of the 1-form: $A^{(1)}=\frac{1}{2!}\,\big{[}\varepsilon_{\mu\nu\sigma}\,A^{\sigma}\big{]}\,(dx^{\mu}\wedge dx^{\nu})\sim\frac{1}{2!}\,B_{\mu\nu}\,(dx^{\mu}\wedge dx^{\nu})$ is the 2-form. These observations are the reasons behind the numeral factors that appear in the discrete duality transformations (16) modulo the overall factors of $\pm\,i$.:
@@
where (i) the gauge-fixing term of the Abelian 1-form field interchanges with the kinetic term of the Abelian 2-form field, and (ii) the gauge-fixing term of the Abelian 2-form field interchanges with the kinetic term of the Abelian 1-form field. This observation is very crucial as it will play a key role at the quantum level when we shall discuss various kinds of continuous and discrete symmetries of our $3D$ field-theoretic system (as the BRST-quantized theory).
@@ 3 BRST and co-BRST Invariant Lagrangian Density
The gauge-fixed Lagrangian density (12) for our $3D$ combined system of the free Abelian 1-form and 2-form gauge theories can be linearized, in its most general form in two different ways by invoking the auxiliary fields, as follows
${\cal L}^{(1)}_{({\cal B},B)}$ $=$ $\dfrac{1}{2}\,{\cal B}^{\mu}{\cal B}_{\mu}-{\cal B}^{\mu}\Big{(}\varepsilon_{\mu\nu\sigma}\partial^{\nu}A^{\sigma}-\dfrac{1}{2}\,\partial_{\mu}\widetilde{\phi}\Big{)}-B\,(\partial\cdot A)+\dfrac{B^{2}}{2}$ (18)
@@
where the subscripts on the above Lagrangian densities denote the short forms of the Nakanishi-Lautrup auxiliary fields that have been invoked to linearize the kinetic and gauge-fixing terms of the combined system of the free Abelian 2-form and 1-form gauge theories. To be more specific, we note that, in the Lagrangian density (18), a pair of auxiliary fields (${\cal B}_{\mu},\,B$) have been invoked to linearize the kinetic and gauge-fixing terms of the $3D$ Abelian 1-form gauge theory, respectively. On the other hand, another pair of auxiliary fields (${\cal B},\,B_{\mu}$) have been incorporated into the theory to linearize the kinetic and gauge-fixing terms of the $3D$ Abelian 2-form theory, respectively. In exactly similar fashion, we can specify the concise form of the subscript on the Lagrangian density (19) where the replacements for the pairs of auxiliary fields are: $({\cal B}_{\mu},\,B)\Rightarrow(\bar{\cal B}_{\mu},\,B)$ and $({\cal B},\,B_{\mu})\Rightarrow({\cal B},\,\bar{B}_{\mu})$, respectively. In the above Lagrangian densities (18) and (19), an additional pair of dynamical fields ($\widetilde{\phi},\;\phi$) have been added at appropriate places with proper mass dimensions in natural units (where $\hbar=c=1$). These fields are nothing but the pseudo-scalar and scalar fields, respectively (cf. Sec. 5 for more discussions). The above linearizations are general in the sense that we have taken into account both the signs that can be associated with the fields ($\widetilde{\phi},\;\phi$). In other words, a close look at (18) and (19) demonstrates that both the Lagrangian densities can be exchanged with each-other by taking into account the replacements: $(\widetilde{\phi},\;\phi)\leftrightarrow(-\,\widetilde{\phi},\;-\,\phi),\;({\cal B}_{\mu},\,B)\leftrightarrow(\bar{\cal B}_{\mu},\,B),\;({\cal B},\,B_{\mu})\leftrightarrow({\cal B},\,\bar{B}_{\mu})$. At this juncture, it is interesting to point out that the discrete duality symmetry transformations (16), existing for the gauge-fixed Lagrangian density (12), can be generalized into the following transformations:
@@
For the sake of brevity, in our present endeavor, we consider only (i) the off-shell nilpotent (co-)BRST symmetry transformations which are the generalizations of the classical (dual-)gauge symmetry transformations (14) and (2) to the quantum level, and (ii) the (co-)BRST invariant Lagrangian density ${\cal L}_{B}$ which is the generalization of the classical linearized version of the Lagrangian density ${\cal L}^{(1)}_{({\cal B},B)}$ [cf. Eq. (18)] to its quantum level. In other words, we promote the Lagrangian density (18) to incorporate into itself the Faddeev-Popov (FP) ghost terms corresponding to the $3D$ field-theoretic model of the combined system of the free Abelian 1-form and 2-form gauge theories as follows
@@
where the FP-terms, corresponding to the free massless Abelian 2-form theory, have been systematically derived in our earlier works [6, 9]. In the latter work [9], we have considered the BRST-quantized version of the $4D$ Stückelberg-modified massive Abelian 2-form theory and the FP-ghost term (i.e. $\partial^{\mu}\bar{C}\,\partial_{\mu}C$) for the free Abelian 1-form theory is the standard book material. In the above Lagrangian density (21), the Lorentz vector fermionic (i.e. $C_{\mu}^{2}=\bar{C}_{\mu}^{2}=0,\,C_{\mu}\,\bar{C}_{\nu}+\bar{C}_{\nu}\,C_{\mu}=0,\,C_{\mu}\,C_{\nu}+C_{\nu}\,C_{\mu}=0,$ etc.) (anti-)ghost fields $(\bar{C}_{\mu})C_{\mu}$ are the generalizations of the gauge transformation parameter $\Lambda_{\mu}$ [cf. Eq. (2)] and they carry the ghost numbers (1)+1(-1)+1, respectively, and the bosonic (anti-)ghost fields $(\bar{\beta})\beta$ are endowed with the ghost numbers (2)+2(-2)+2. Another pair of the fermionic (i.e. $C^{2}=\bar{C}^{2}=0,\,C\,\bar{C}+\bar{C}\,C=0$) (anti-)ghost fields $(\bar{C})C$ carry the ghost numbers (1)+1(-1)+1 and they are the generalizations of the Lorentz scalar gauge transformation parameter $\Lambda$ [cf. Eq. (2)] and the auxiliary (anti-)ghost fields $(\rho)\lambda$ are endowed with ghost numbers (1)+1(-1)+1, respectively. These (anti-)ghost fields are required to maintain the sacrosanct property of unitarity at any arbitrary order of perturbative computations. The other symbols in (21) carry the same meaning as explained in the context of the Lagrangian density in (18).
@@
$C_{\mu}\,\longrightarrow\,\pm\,i\,\bar{C}_{\mu},\qquad\bar{C}_{\mu}\,\longrightarrow\,\pm\,i\,C_{\mu},\qquad C\,\longrightarrow\,\mp\,i\,\bar{C},\qquad\bar{C}\,\longrightarrow\,\mp\,i\,C,$
$\beta\,\longrightarrow\,\pm\,i\,\bar{\beta},\qquad\quad\bar{\beta}\,\longrightarrow\,\mp\,i\,\beta,\qquad\;\;\;\;\rho\,\longrightarrow\,\mp\,i\,\lambda,~\qquad\lambda\,\longrightarrow\,\mp\,i\,\rho.$ (25)
Thus, it is clear that the total (co-)BRST invariant Lagrangian density (21) respects the discrete duality symmetry transformations in the non-ghost sector [cf. Eq. (3)] as well as in the ghost-sector [cf. Eq. (3)]. To sum-up, the Lagrangian density (21) respects the infinitesimal, continuous and nilpotent symmetries [cf. Eqs. (3), (3)] and a set of discrete duality symmetry transformations. The latter is the sum of the discrete symmetry transformations that have been listed in (3) and (3). We call these discrete symmetry transformations as the duality symmetry transformations because the basic 1-form and 2-form fields are connected with each-other [cf. Eq. (3)] and they are dual to each-other on a flat $3D$ Minkowskian spacetime manifold [cf. footnote below Eq. (16)].
@@ 4 Bosonic Symmetry and Algebraic Structures
It is straightforward to note, from the off-shell nilpotent symmetries in our equations (3) and (3), that the anticommutator (i.e. $\{s_{b},\;s_{d}\}$) between them is not equal to zero. In fact, this anticommutator defines a set of the bosonic symmetry (i.e. $s_{\omega}=\{s_{b},\;s_{d}\}$) transformations, under which, the fields of (21) transform as:
$s_{\omega}B_{\mu\nu}=\varepsilon_{\mu\nu\sigma}\,\partial^{\sigma}B+\big{(}\partial_{\mu}{\cal B}_{\nu}-\partial_{\nu}{\cal B}_{\mu}\big{)},\qquad\quad s_{\omega}A_{\mu}=-\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}B^{\sigma}+\partial_{\mu}{\cal B},$
$s_{\omega}C_{\mu}=\partial_{\mu}\lambda,\quad s_{\omega}\bar{C}_{\mu}=\partial_{\mu}\rho,\quad s_{\omega}\big{[}B_{\mu},\,{\cal B}_{\mu},\,B,\,{\cal B},\,\phi,\,\widetilde{\phi},\,\beta,\,\bar{\beta},\rho,\,\lambda,\,C,\,\bar{C}\big{]}=0.$ (26)
At this stage, it is worthwhile to mention that under the above bosonic symmetry transformations, the (anti-)ghost fields either do not transform at all or they transform up to the $U(1)$ gauge symmetry-type transformations. Furthermore, it is very interesting to point out that, in their operator forms, the nilpotent (co-)BRST symmetry transformations and the bosonic symmetry transformation obey the following algebra, namely;
@@
which establish the fact that the bosonic symmetry transformation, in its operator form, commutes with both the off-shell nilpotent (co-)BRST symmetry transformation operators. The algebra (4) is reminiscent of the following algebra obeyed by a set of three de Rham cohomological operators of differential geometry [14, 15, 16, 17, 18]
@@
where $d$ (with $d^{2}=0$) is the exterior derivative, $\delta=\pm\,*\,d\,*$ (with $\delta^{2}=0$) is the co-exterior (or dual-exterior) derivative and $\Delta=(d+\delta)^{2}$ is the Laplacian operator (which is always positive definite). Here the symbol * stands for the Hodge duality operator on the compact spacetime manifold on which the cohomological operators are defined. The above algebraic structure, satisfied by the de Rham cohomological operators, are popularly known as the Hodge algebra of differential geometry (see, e.g. [14, 15, 16, 17, 18] for details).
We end this short section with a couple of useful remarks. First of all, at the algebraic level, it is very tempting to appreciate the similarity between the above algebraic relationships in (4) and (4) which allow us to state that there is a one-to-one mapping******We have obtained the one-to-one mapping because we have considered only the Lagrangian density (21) which is the generalization of the linearized classical version of the gauge-fixed Lagrangian density (18) to its quantum level. If we had considered the quantum version of the linearized classical version of the gauge-fixed Lagrangian density (19) together with the above generalization, we would have ended up with the two-to-one mapping between the symmetry transformation operators and the de Rham cohomological operators of differential geometry as we have obtained in our earlier works [6, 7, 8, 9]. between the symmetry transformation operators ($s_{b},\;s_{d},\;s_{\omega}$) and the de Rham cohomological operators ($d,\;\delta,\;\Delta$). To corroborate this claim, we have to recall our key observations in Secs. 2 and 3 where we have established that, under the (dual-)gauge and (co-)BRST symmetry transformations, the (i) gauge-fixing terms (owing their origin to the co-exterior derivative), and (ii) kinetic terms (owing their existence to the exterior derivative) for the gauge fields remain invariant, respectively. Hence, the above identification: $(s_{b},\;s_{d},\;s_{\omega})\Leftrightarrow(d,\;\delta,\;\Delta)$ is realistic at the algebraic level. Second, the (co-)exterior derivatives $(\delta)d$, as pointed out earlier, are connected by the relationship: $\delta=\pm\,*\,d\,*$. Within the framework of BRST formalism, this relationship is realized in terms of the nilpotent (co-)BRST transformation operators and the discrete transformation operators [cf. Eqs. (3), (3)] as
@@
where the symbol * stands for the discrete duality symmetry transformations. Thus, we observe that the interplay between the continuous and discrete symmetry operators of our theory provides the physical realization of the relationship: $\delta=\pm\,*\,d\,*$. In the above equation (29), as is obvious, the generic field of the Lagrangian density (21) has been denoted by the field $\Phi$. The ($\pm$) signs, on the r.h.s. of the above equation (29), are dictated by a couple of successive operations of the discrete duality symmetry transformation operator on the generic field $\Phi$ of the Lagrangian density (21) as follows [31]:
@@
It is an elementary exercise to note that, for all the bosonic fields of our (co-)BRST invariant Lagrangian density (21), we have the plus sign on the r.h.s. of (30) and, hence, in the equation (29), too. For instance, it is straightforward to check that: $(*\,B_{\mu\nu})=+\,B_{\mu\nu}$. As a consequence, we observe that: $s_{d}\,B_{\mu\nu}=+\,*\,s_{b}\,*\,B_{\mu\nu}$ is true. On the other hand, the signs on the r.h.s. of (30) and (29) are negative when we take into account the fermionic fields of the (co-)BRST invariant Lagrangian density (21). For instance, we can readily check that: $(*\,C_{\mu})=-\,C_{\mu}$. As a consequence, we have the algebraic relationship between the off-shell nilpotent (co-)BRST symmetry transformation operators [cf. Eqs. (3), (3)] as: $s_{d}\,C_{\mu}=-\,*\,s_{b}\,*\,C_{\mu}$ where the symbol * stands for the set of discrete duality symmetry transformation operators that have been listed in the equations (3) and (3) for our (co-)BRST invariant Lagrangian density (21). To be precise, these discrete duality symmetry transformations [cf. Eqs. (3), (3)] are present in the non-ghost and ghost-sectors of our field-theoretic model of the combined system of the BRST-quantized version of the $3D$ free Abelian 1-from and 2-form gauge theories.